OpenStudy (anonymous):

The Better Baby Buggy Co. has just come out with a new model, the Turbo. The market research department predicts that the demand equation for Turbos is given by q = −2p + 288, where q is the number of buggies it can sell in a month if the price is $p per buggy. At what price should it sell the buggies to get the largest revenue? 1.$=__________ 2.What is the largest monthly revenue? $___________

OpenStudy (anonymous):

giving medals out

OpenStudy (anonymous):

Revenue R = Quantity Q * Unit Price P = (−2P + 288) * P = - 2P^2 + 288P Maximize Revenue: R' = -4P + 288 R' = 0 --> P = 288/4 = $72

OpenStudy (anonymous):

Largest monthly revenue: R( 72) = - 2 (72)^2 + 288 * 72 = $10,368